### v3

# Improving Graph Neural Network Training Efficiency By Using Top Non-Robust Samples In The Training Set

Abstract
Graph Neural Networks (GNNs) are a highly effective neural network architecture for processing graph-structured data. Unlike traditional neural networks that rely solely on the features of the data as input, GNNs leverage both the graph structure, which represents the relationships between data points, and the feature matrix of the data to optimize their feature representation. This unique capability enables GNNs to achieve superior performance across various tasks. However, it also makes GNNs more susceptible to noise from both the graph structure and the data features, which can significantly degrade their performance in common tasks such as classification and prediction. To address this issue, this paper proposes a novel method for constructing training sets by identifying training samples that are particularly sensitive to noise for a given model. These samples are then used to enhance the model’s ability to handle noise-prone instances effectively. Experimental results demonstrate that this approach can significantly improve training efficiency.
Index Terms:
Graph Neural Networks, Training, Efficiency
I Introduction
The rise of Graph Neural Networks (GNNs) has brought about a transformative shift in the field of machine learning, particularly for tasks involving graph-structured data [1, 2, 3, 4]. These networks have demonstrated success across a wide array of real-world scenarios, including recommendation systems [5], traffic forecasting [6], chip design [7], and social network analytics[8]. The unique advantage of GNNs lies in their ability to effectively leverage the graph that represents the relationships between data points to optimize their feature representations. This enables GNNs to fully utilize the crucial graph topology inherent in graph-structured data, unlike traditional neural networks, which optimize data representations based solely on feature vectors, thus overlooking the valuable graph structure.
However, like other neural networks, GNNs are susceptible to noise in the feature space[9, 10]. Additionally, the introduction of graph structure introduces another layer of interference, as GNNs are also affected by noise within the graph itself. Research has shown that even slight modifications to the graph—such as adding, removing, or rearranging edges—can significantly impact the performance of GNNs [11, 12]. The feature vectors contaminated by noise, along with erroneous and redundant edges in the graph topology, can be viewed as adversarial attacks that mislead the GNNs. To address this issue, researchers have proposed various methods to reduce the impact of noise on the model’s performance. These approaches can be divided into two categories: reactive and proactive defenses. Reactive defenses focus on identifying adversarial examples within the model’s inputs, as explored in previous studies[13, 14, 15]. On the other hand, proactive defenses aim to strengthen the model’s robustness, making it less susceptible to adversarial attacks, as demonstrated by methods proposed in works such as [16, 17].
However, due to the diversity of noise types and model structures, it is challenging to develop effective methods for defending against noise across different scenarios. Unlike existing approaches that focus on mitigating the effects of noise, we argue that, given the inevitability of noise, the focus should be on how to better leverage it. In this paper, we propose a method that first identifies the training samples most vulnerable to noise from the dataset, and then uses these samples to construct a training set for training the GNN model. This approach effectively enhances the model’s ability to process noisy data correctly. Our experiments on multiple benchmark datasets demonstrate the effectiveness of this method.
II Preliminaries
II-A The Vulnerability of GNNs to Noise
The vulnerability of Graph Neural Networks (GNNs) to noise refers to the difficulty GNNs face in maintaining consistent outputs when exposed to perturbations in the graph structure, such as the presence of incorrect or redundant edges in the input graph, or errors in the feature representations within the input feature matrix. [20] investigated the task- and model-agnostic vulnerabilities of GNNs, showing that these networks are very vulnerable to adversarial perturbations, regardless of the downstream task.
II-B Spectral Graph Theory
Spectral Graph Theory is a mathematical method that uses the eigenvalues and eigenvectors of the Laplacian matrix of a graph to study its properties. Given a graph $G=(V,E,w)$, where $V$ represents the set of vertices, $E$ is the set of edges, and $w$ is a function assigning positive weights to each edge. The Laplacian matrix of graph $G$, which is a symmetric diagonally dominant (SDD) matrix, is defined as follows:
$L(p,q)=\begin{cases}-w(p,q)&\text{if }(p,q)\in E,\\ \sum\limits_{(p,t)\in E}w(p,t)&\text{if }p=q,\\ 0&\text{otherwise}.\end{cases}$ (1)
For better numerical stability and to make the Laplacian matrix scale-invariant, the normalized Laplacian matrix $L_{\text{norm}}$ can be computed as:
$L_{\text{norm}}=D^{-1/2}LD^{-1/2}$
According to spectral graph theory, when analyzing a graph, often only one or a few special eigenvalues and their corresponding eigenvectors are useful. For example, when analyzing the clustering structure of a graph using spectral graph theory, only the smallest eigenvalues and their corresponding eigenvectors are useful[18]. Therefore, for a Laplacian matrix, it is unnecessary to compute all the eigenpairs through direct eigen decomposition, and for a large-scale graph, the computational cost would be too high. The Courant-Fischer Minimax Theorem[19] allows the eigenvalues to be computed iteratively without explicitly calculating all the eigenvalues. Specifically:
The k-th largest eigenvalue of the Laplacian matrix $L\in\mathbb{R}^{|V|\times|V|}$ can be computed as follows:
$\lambda_{\textit{k}}(L)=\min_{\dim(U)=k}\max_{x\in U,\,x\neq 0}\frac{x^{T}Lx}{x^{T}x}$
In many practical applications, the problem-solving process involves not just a single matrix, but rather the relationship between two matrices. Therefore, the Courant-Fischer Minimax Theorem is extended to the generalized case, resulting in the following Generalized Courant-Fischer Minimax Theorem:
Consider two Laplacian matrices $L_{X}\in\mathbb{R}^{|V|\times|V|}$ and $L_{Y}\in\mathbb{R}^{|V|\times|V|}$, where it holds that $\text{null}(L_{Y})\subseteq\text{null}(L_{X})$. Under the condition $1\leq k\leq\text{rank}(L_{Y})$, the $k$-th largest eigenvalue of the matrix $L_{Y}^{+}L_{X}$ can be computed as follows:
$\lambda_{k}(L_{Y}^{+}L_{X})=\min_{\dim(U)=k,\ U\perp\text{null}(L_{Y})}\max_{x\in U}\frac{x^{T}L_{X}x}{x^{T}L_{Y}x}$
III Method
Figure 1: Overview of the proposed method.
Figure 1 shows the proposed method for training GNNs using non-robust training samples, which consists of four key phases: Phase (1) is GNN embedding, which uses the original input to train a GNN and takes the output of the hidden layer as the embedding result of the data’s feature matrix; In Phase (2), we use k-NN graphs to capture the underlying manifolds of the input feature matrix and the embedded feature matrix; In Phase (3), we calculate each node’s level of non-robustness using the input manifold and the output manifold; In Phase (4), we use the top non-robust samples selected from the training set to train the GNN. The details of these four phases are elaborated in the following sections.
III-A Phase 1: GNN Embedding
[21] proposed leveraging a bijective distance mapping between the manifolds of a model’s input and output data to analyze the adversarial robustness of each data point for a given model.
To obtain the output data of the GNN model, we first train the GNN using the original feature matrix and input graph. Once trained, the GNN is then used to transform the input feature matrix. Since the final layer of a GNN is typically task-specific—such as a softmax layer for multi-class classification tasks—the outputs of the hidden layers prior to the final layer are used as embeddings of the original feature
III-B Phase 2: Manifold Capturing
The manifold of the data is typically captured using graphs, and k-nearest neighbor graphs are the most widely used for this purpose. In this phase, we use k-NN graphs to capture the underlying manifolds of the input feature matrix and the embedded feature matrix. The conventional kNN graph construction algorithm exhibits a time complexity of $O(|N|^{2})$, where $|N|$ is the total number of nodes in the graph. Such quadratic complexity makes it impractical for large-scale datasets. To overcome this limitation, an approximate kNN graph construction algorithm can be employed. This approach leverages an extension of the probabilistic skip list structure to achieve a reduced time complexity of $O(|N|\log|N|)$, which is more suitable for large-scale data processing [22].
III-C Phase 3: Adversarial Robustness Evaluation
Building on the theoretical framework introduced in [21], we leverage the largest generalized eigenvalues and their associated eigenvectors of $L_{Y}^{+}L_{X}$ to assess the stability of individual nodes in GNNs. Specifically, we calculate the weighted eigensubspace matrix $V_{s}\in\mathbb{R}^{|V|\times s}$ to perform spectral embedding of the input manifold $G_{X}=(V,E_{X})$, where $|V|$ represents the total number of nodes. The matrix $V_{s}$ is expressed as:
$V_{s}=\begin{bmatrix}v_{1}\sqrt{\zeta_{1}},&v_{2}\sqrt{\zeta_{2}},&\ldots,&v_{s}\sqrt{\zeta_{s}}\end{bmatrix},$
where $\zeta_{1}\geq\zeta_{2}\geq\cdots\geq\zeta_{s}$ are the top $s$ eigenvalues of $L_{Y}^{+}L_{X}$, and $v_{1},v_{2},\ldots,v_{s}$ are the corresponding eigenvectors.
Using $V_{s}$, we project the nodes of $G_{X}$ into an $s$-dimensional space, representing each node $p$ as the $p$-th row of $V_{s}$. The stability of an edge $(p,q)\in E_{X}$ is then quantified by computing the spectral embedding distance between nodes $p$ and $q$:
$\|V_{s}^{\top}e_{p,q}\|_{2}^{2}.$
A metric, referred to as the Spade score, is defined as the average embedding distance to the neighbors within the input manifold to assess the robustness of a node $p$:
$\text{Spade}(p)=\frac{1}{|N_{X}(p)|}\sum_{q\in N_{X}(p)}\|V_{s}^{\top}e_{p,q}\|_{2}^{2},$
where $N_{X}(p)$ represents the set of neighbors of node $p$ in $G_{X}$. A higher $\text{Spade}(i)$ value suggests that node $i$ is more prone to adversarial attacks.
III-D Phase 4: GNN Training
For the samples in the training set, we rank them in descending order based on their Spade scores and select the top portion to form a new training set. The points in this new training set are those that are highly susceptible to noise and likely to mislead the GNN. By training the GNN using only these noise-sensitive samples, we are effectively enhancing the GNN’s ability to resist being misled by such samples.
IV Experiments
IV-A Experimental Setup
We use three standard citation network benchmark datasets to evaluate the proposed method: Cora, Citeseer, and Pubmed. In all these citation networks, nodes represent documents and edges correspond to citations. Each node is associated with a sparse bag-of-words feature vector and a class label.
The Cora dataset consists of 2,708 scientific publications categorized into seven classes, forming a citation network with 5,429 links.
The Citeseer dataset includes 3,327 scientific publications categorized into six classes, with a total of 4,732 links in the citation network. Compared to Cora, Citeseer is sparser and exhibits a more uneven class distribution, making it more challenging for graph-based classification methods.
The Pubmed dataset contains 19,717 scientific publications from the field of biomedicine, categorized into three classes. Each publication is represented by a Term Frequency-Inverse Document Frequency (TF-IDF) weighted word vector derived from a vocabulary of 500 words. The citation network in Pubmed includes 44,338 links, making it significantly larger than both Cora and Citeseer.
We extract 20% of the samples from the standard Cora dataset as the test set. From the remaining 80% of the samples, we construct two training sets: one by randomly selecting 80% of the samples and the other by selecting the top 80% non-robust nodes using our method. Both training sets are used to train GCNs with the same architecture and training parameters. To further ensure a fair comparison, we initialize both GCNs with the same random seed by setting torch.manual_seed(42), which guarantees identical random initialization of their parameters.
IV-B Experimental Results
IV-B1 Experimental Results on Graph Convolutional Networks (GCNs)
For the Cora dataset, the classification accuracy results on the test set as training epochs progress is shown in Figure 2 and Table I .
TABLE I: Classification Accuracy Across Epochs Using GCN (Cora)
Number of Epochs 1 2 3 4 5 6
Random Training Set 0.3142 0.3475 0.4140 0.4880 0.6081 0.7024
Our Training Set 0.3826 0.4991 0.5490 0.6839 0.7043 0.8004
Figure 2: Comparison of Classification Accuracy Across Epochs Using GCN (Cora).
For the Citeseer dataset, the classification accuracy results on the test set as training epochs progress is shown in Figure 3 and Table II .
TABLE II: Classification Accuracy Across Epochs Using GCN (Citeseer)
Number of Epochs 1 2 3 4 5 6
Random Training Set 0.2030 0.4120 0.5158 0.5774 0.6120 0.6677
Our Training Set 0.2286 0.4872 0.5083 0.5865 0.6511 0.7368
Figure 3: Comparison of Classification Accuracy Across Epochs Using GCN (Citeseer).
For the Pubmed dataset, the classification accuracy results on the test set as training epochs progress is shown in Figure 4 and Table III .
TABLE III: Classification Accuracy Across Epochs Using GCN (Pubmed)
Number of Epochs 1 2 3 4 5 6
Random Training Set 0.3918 0.3974 0.5044 0.5970 0.6163 0.6208
Our Training Set 0.3974 0.6224 0.6348 0.6515 0.6878 0.6774
Figure 4: Comparison of Classification Accuracy Across Epochs Using GCN (Pubmed).
IV-B2 Experimental Results on Graph Attention Networks (GATs)
For the Cora dataset, the classification accuracy results on the test set as training epochs progress is shown in Figure 5 and Table IV .
TABLE IV: Classification Accuracy Across Epochs Using GAT (Cora)
Number of Epochs 1 2 3 4 5 6
Random Training Set 0.3124 0.3623 0.4603 0.6488 0.7634 0.8022
Our Training Set 0.4713 0.6617 0.7338 0.8022 0.8373 0.8466
Figure 5: Comparison of Classification Accuracy Across Epochs Using GAT (Cora).