### v9

# Improving Graph Neural Network Training, Defense and Hypergraph Partitioning via Adversarial Robustness Evaluation

Abstract
Graph Neural Networks (GNNs) are a highly effective neural network architecture for processing graph-structured data. Unlike traditional neural networks that rely solely on the features of the data as input, GNNs leverage both the graph structure, which represents the relationships between data points, and the feature matrix of the data to optimize their feature representation. This unique capability enables GNNs to achieve superior performance across various tasks. However, it also makes GNNs more susceptible to noise and adversarial attacks from both the graph structure and data features, which can significantly increase the training difficulty and degrade their performance. Similarly, a hypergraph is a highly complex structure, and partitioning a hypergraph is a challenging task. This paper leverages spectral adversarial robustness evaluation to effectively address key challenges in complex-graph algorithms. By using spectral adversarial robustness evaluation to distinguish robust nodes from non-robust ones and treating them differently, we propose a training-set construction strategy that improves the training quality of GNNs. In addition, we develop algorithms to enhance both the adversarial robustness of GNNs and the performance of hypergraph partitioning. Experimental results show that this series of methods is highly effective.
footnotetext: Correspondence to: yongyuw@mtu.edu
1 Introduction
The rise of Graph Neural Networks (GNNs) has brought about a transformative shift in the field of machine learning, particularly for tasks involving graph-structured data [1, 2, 3, 4]. These networks have demonstrated success across a wide array of real-world scenarios, including recommendation systems [5], traffic forecasting [6], chip design [7], language processing [8] and social network analytics[9]. The unique advantage of GNNs lies in their ability to effectively leverage the graph that represents the relationships between data points to optimize their feature representations. This enables GNNs to fully utilize the crucial graph topology inherent in graph-structured data, unlike traditional neural networks, which optimize data representations based solely on feature vectors, thus overlooking the valuable graph structure.
However, like other neural networks, GNNs are susceptible to noise and adversarial attacks in the graph and feature space[10, 11]. . Research has shown that even slight modifications to the graph—such as adding, removing, or rearranging a small number of edges—can significantly impact the performance of GNNs [12, 13].
To address this issue, researchers have proposed various methods to reduce the impact of noise on the model’s performance. These approaches can be divided into two categories: reactive and proactive defenses. Reactive defenses focus on identifying adversarial examples within the model’s inputs, as explored in previous studies[14, 15, 16]. On the other hand, proactive defenses aim to strengthen the model’s robustness, making it less susceptible to adversarial attacks, as demonstrated by methods proposed in works such as [17, 18].
However, due to the diversity of noise types and model structures, developing effective methods to defend against noise across various scenarios remains a significant challenge. Unlike existing approaches that focus on mitigating the effects of noise, we argue that, given the inevitability of noise, the focus should shift toward how to better leverage it. From this new perspective, this paper proposes a framework that first identifies the most noise-sensitive training samples and then uses these samples to construct a smaller but more effective training set for training the GNN model. This approach enhances the model’s ability to correctly process noisy data.
In this work, we also propose a new approach to improve the defense ability of GNNs. Specifically, we apply spectral adversarial robustness evaluation to partition nodes in the graph into two groups: robust nodes that are relatively insensitive to noise and perturbations, and non-robust nodes that are more easily affected. We then treat these two groups differently. For robust nodes, we induce a subgraph containing only them from the original graph, and use a GNN to predict their class labels. For non-robust nodes, we leverage the topology of the original graph to propagate class information from robust nodes to obtain their labels
We also provide a efficient high performance hypergraph partitioning method based on spectral adversarial robustness evaluation and demonstrate its effectiveness on benchmark VLSI data sets.
2 Preliminaries
2.1 The Vulnerability of GNNs to noise and adversarial attacks
The vulnerability of Graph Neural Networks (GNNs) to noise and adversarial attacks refers to the difficulty GNNs face in maintaining consistent outputs when exposed to perturbations in the graph structure, such as the presence of incorrect or redundant edges in the input graph, or errors in the feature representations within the input feature matrix. [23] investigated the task- and model-agnostic vulnerabilities of GNNs, showing that these networks are very vulnerable to adversarial perturbations, regardless of the downstream task.
2.2 Spectral Graph Theory
Spectral Graph Theory is a mathematical method that uses the eigenvalues and eigenvectors of the Laplacian matrix of a graph to study its properties. Given a graph $G=(V,E,w)$, where $V$ represents the set of vertices, $E$ is the set of edges, and $w$ is a function assigning positive weights to each edge. The Laplacian matrix of graph $G$, which is a symmetric diagonally dominant (SDD) matrix, is defined as follows:
$L(p,q)=\begin{cases}-w(p,q)&\text{if }(p,q)\in E,\\ \sum\limits_{(p,t)\in E}w(p,t)&\text{if }p=q,\\ 0&\text{otherwise}.\end{cases}$ (1)
For better numerical stability and to make the Laplacian matrix scale-invariant, the normalized Laplacian matrix $L_{\text{norm}}$ can be computed as:
$L_{\text{norm}}=D^{-1/2}LD^{-1/2}$
According to spectral graph theory, analyzing a graph often requires only one or a few specific eigenvalues and their corresponding eigenvectors [20]. For example, when analyzing the clustering structure of a graph using spectral graph theory, only the smallest eigenvalues and their corresponding eigenvectors are useful[21]. Therefore, for a Laplacian matrix, it is unnecessary to compute all the eigenpairs through direct eigen decomposition, and for a large-scale graph, the computational cost would be too high. The Courant-Fischer Minimax Theorem[22] allows the eigenvalues to be computed iteratively without explicitly calculating all the eigenvalues. Specifically:
The k-th largest eigenvalue of the Laplacian matrix $L\in\mathbb{R}^{|V|\times|V|}$ can be computed as follows:
$\lambda_{\textit{k}}(L)=\min_{\dim(U)=k}\max_{x\in U,\,x\neq 0}\frac{x^{T}Lx}{x^{T}x}$
In many practical applications, the problem-solving process involves not just a single matrix, but rather the relationship between two matrices. Therefore, the Courant-Fischer Minimax Theorem is extended to the generalized case, resulting in the following Generalized Courant-Fischer Minimax Theorem:
Consider two Laplacian matrices $L_{X}\in\mathbb{R}^{|V|\times|V|}$ and $L_{Y}\in\mathbb{R}^{|V|\times|V|}$, where it holds that $\text{null}(L_{Y})\subseteq\text{null}(L_{X})$. Under the condition $1\leq k\leq\text{rank}(L_{Y})$, the $k$-th largest eigenvalue of the matrix $L_{Y}^{+}L_{X}$ can be computed as follows:
$\lambda_{k}(L_{Y}^{+}L_{X})=\min_{\dim(U)=k,\ U\perp\text{null}(L_{Y})}\max_{x\in U}\frac{x^{T}L_{X}x}{x^{T}L_{Y}x}$
3 Improving Graph Neural Network Training
3.1 Method
Figure 1: Overview of the proposed method.
Figure 1 shows the proposed method for training GNNs using top non-robust training samples in the training set, which consists of four key phases: Phase (1) is GNN embedding, which uses the original input to train a GNN and takes the output of the hidden layer as the embedding result of the data’s feature matrix; In Phase (2), we use k-NN graphs to capture the underlying manifolds of the input feature matrix and the embedded feature matrix; In Phase (3), we calculate each node’s level of non-robustness using the input manifold and the output manifold; In Phase (4), we use the top non-robust samples selected from the training set to train the GNN. In the rest of this section, we provide a detailed description of these four phases.
3.1.1 Phase 1: GNN Embedding
[24] proposed leveraging a bijective distance mapping between the manifolds of a model’s input and output data to analyze the adversarial robustness of each data point for a given model.
To obtain the output data of the GNN model, we first train the GNN using the original feature matrix and input graph. Once trained, the GNN is then used to transform the input feature matrix. Since the final layer of a GNN is typically task-specific—such as a softmax layer for multi-class classification tasks—we take the output of the hidden layer before the final output layer as the embedding result.
3.1.2 Phase 2: Manifold Capturing
The manifold of the data is typically captured using graphs, and k-nearest neighbor graphs are the most widely used for this purpose. In this phase, we use k-nearest neighbor (k-NN) graphs to capture the underlying manifolds of the input feature matrix and the embedded feature matrix. The conventional k-NN graph construction algorithm exhibits a time complexity of $O(|N|^{2})$, where $|N|$ is the total number of nodes in the graph. Such quadratic complexity makes it impractical for large-scale datasets. To overcome this limitation, an approximate kNN graph construction algorithm can be employed [25]. This approach leverages an extension of the probabilistic skip list structure to achieve a reduced time complexity of $O(|N|\log|N|)$, which is more suitable for large-scale data processing.
3.1.3 Phase 3: Adversarial Robustness Evaluation
Let $L_{X}$ and $L_{Y}$ denote the Laplacian matrices corresponding to the k-NN graphs of the input feature matrix and the embedded feature matrix constructed in phase 2, respectively. Building on the theoretical framework introduced in [24], we leverage the largest generalized eigenvalues and their associated eigenvectors of $L_{Y}^{+}L_{X}$ to assess the robustness of each individual sample to noise during the processing by a given model. Specifically, we calculate the weighted eigensubspace matrix $V_{s}\in\mathbb{R}^{|V|\times s}$ to perform spectral embedding of the input manifold $G_{X}=(V,E_{X})$, where $|V|$ represents the total number of nodes. The matrix $V_{s}$ is expressed as:
$V_{s}=\begin{bmatrix}v_{1}\sqrt{\zeta_{1}},&v_{2}\sqrt{\zeta_{2}},&\ldots,&v_{s}\sqrt{\zeta_{s}}\end{bmatrix}$,
where $\zeta_{1}\geq\zeta_{2}\geq\cdots\geq\zeta_{s}$ are the top $s$ eigenvalues of $L_{Y}^{+}L_{X}$, and $v_{1},v_{2},\ldots,v_{s}$ are the corresponding eigenvectors.
Using $V_{s}$, we project the nodes of $G_{X}$ into an $s$-dimensional space, representing each node $p$ as the $p$-th row of $V_{s}$. According to [24], the stability of an edge $(p,q)\in E_{X}$ is quantified by computing the spectral embedding distance between nodes $p$ and $q$:
$\|V_{s}^{\top}e_{p,q}\|_{2}^{2}.$
To assess the robustness of a node $p$, a metric referred to as the Spade score is defined as the average embedding distance to the neighbors within the input manifold:
$\text{Spade}(p)=\frac{1}{|N_{X}(p)|}\sum_{q\in N_{X}(p)}\|V_{s}^{\top}e_{p,q}\|_{2}^{2},$
where $N_{X}(p)$ represents the set of neighbors of node $p$ in $G_{X}$. A higher $\text{Spade}(p)$ value suggests that node $p$ is more prone to adversarial attacks.
3.1.4 Phase 4: GNN Training
For the samples in the training set, we rank them in descending order based on their Spade scores and select the top portion to form a new, smaller training set. The points in this new training set are those that are highly susceptible to noise and likely to mislead the GNN. By training the GNN using only these noise-sensitive samples, we are effectively enhancing the GNN’s ability to resist being misled by such samples.
3.2 Experiments
We have evaluated our framework on three of the most popular GNN models and three classic benchmark datasets.
3.2.1 Data Sets
We evaluate the proposed method on three standard citation network benchmarks, where nodes represent documents, edges correspond to citations, and each node is associated with a sparse bag-of-words feature vector and a class label. The key statistics of these datasets are summarized in Table 1.
  • Cora: A dataset of scientific publications categorized into seven classes.
  • Citeseer: A dataset of scientific publications that is sparser than Cora and exhibits a more uneven class distribution, making it more challenging for graph-based classification methods.
  • PubMed: A large biomedical dataset where nodes are represented by TF-IDF word vectors from a 500-word vocabulary.
Table 1: Statistics of datasets used in our experiments.
Dataset Nodes Edges Classes Features
Cora 2,708 5,429 7 1,433
Citeseer 3,327 4,732 6 3,703
PubMed 19,717 44,338 3 500
3.2.2 GNN Models
We use the following three commonly used GNN models for our experiments:
  • Graph Convolutional Network (GCN): A widely used model that applies convolution-like operations to graph-structured data by aggregating information from neighboring nodes using a layer-wise propagation rule.
  • Graph Attention Network (GAT): Introduces an attention mechanism to assign different importance weights to neighboring nodes during the message-passing process.
  • GraphSAGE: An inductive model that learns aggregation functions to generate node embeddings by sampling and aggregating features from a fixed number of neighboring nodes.
3.2.3 Detailed Model Implementation and Parameter Settings
  • GCN: We use the same GCN architecture and parameter settings for all three datasets. The model consists of two graph convolutional layers: the first layer transforms node features to a 64-dimensional space, and the second layer maps them to the number of output classes, with ReLU activation applied after the first convolution. For the Cora dataset, the learning rate is set to 0.01. For the Citeseer and PubMed datasets, it is set to 0.1.
  • GAT: For the Cora dataset, the GAT model consists of two layers: the first layer has 4 attention heads with 8 hidden units each, followed by an ELU activation, and the second layer has 1 attention head that outputs the number of class logits. For the Citeseer dataset, the GAT model also has two layers: the first layer uses 8 attention heads with 8 hidden units each, followed by an ELU activation, and the second layer has 1 attention head that outputs the class logits. For the PubMed dataset, the GAT model has the same structure as the Citeseer model but includes layer normalization after each layer and uses a Leaky ReLU activation after the first layer. For the Cora dataset, the learning rate is set to 0.01. For the Citeseer dataset, it is set to 0.1, and for the PubMed dataset, it is set to 0.05.
  • GraphSAGE: The GraphSAGE architecture for the Cora and Citeseer datasets consists of two layers: the first layer uses SAGEConv with 64 hidden units, followed by batch normalization, ReLU activation, and a dropout layer with a probability of 0.5. The second layer uses SAGEConv with 64 input features and outputs the class labels. The architecture for the PubMed dataset also consists of two layers, but the first layer uses SAGEConv with 32 hidden units, followed by batch normalization, ReLU activation, and a dropout layer. The second layer uses SAGEConv with 32 input features and outputs the class labels. For all three datasets, the learning rate is set to 0.05.
For all the experiments, the Adam optimizer is used with a weight decay of $5\times 10^{-4}$, and the loss function for multi-class classification is CrossEntropyLoss. The number of eigenvectors is set to 10, and the value of $k$ in k-NN is set to 20.
3.2.4 Results
We first perform an 80-20 split of the entire dataset by randomly sampling 80% of the samples to form the original full training set and using the remaining 20% as the test set. Then, we construct two different types of training sets: (1) forming the training set by selecting the top 80% of non-robust samples from the original full training set using our method; (2) forming the training set by randomly sampling 80% of the original full training set.
We train the GNN models using the original full training set and the two smaller training sets mentioned above, and the classification accuracy across different epochs is shown in Table 2. In these charts, Random represents the training set constructed by randomly sampling 80% of the original full training set, Ours represents the training set constructed using our method, and Full represents the original full training set.
From the experimental results, it can be observed that in all 9 combinations of the three GNN models and the three datasets, the training set constructed using our method achieved the highest classification accuracy after 5 epochs of training. Moreover, the advantage of our method is particularly evident on the largest dataset, PubMed. On PubMed, our method outperforms the second-best method by 14.1%, 7.4%, and 3.4% in accuracy for the GCN, GAT, and GraphSAGE models, respectively.
Our method not only significantly outperforms the randomly sampled training sets of the same size but also demonstrates substantial advantages compared to the full training sets, which are 20% larger, highlighting the effectiveness of our approach. The experimental results further show that randomly selecting 80% of the training samples from the full training set yields almost no difference in performance compared to using the full set. However, when 80% of the samples are selected using our method, which prioritizes the most non-robust nodes, the training performance is significantly improved. This demonstrates that focusing on non-robust samples during training, rather than uniformly sampling all nodes, enables the model to better adapt to unavoidable noise, resulting in superior performance on the test set.
Table 2: Classification Accuracy Across Epochs for GCN, GAT, and GraphSAGE on Cora, Citeseer, and PubMed
Model Dataset Training Set Epoch 1 Epoch 2 Epoch 3 Epoch 4 Epoch 5
GCN Cora Random 0.2939 0.3290 0.5360 0.7209 0.7671
Ours 0.5416 0.6580 0.6525 0.7375 0.8041
Full 0.2939 0.3290 0.5527 0.7227 0.8041
Citeseer Random 0.2331 0.6030 0.6030 0.6256 0.6256
Ours 0.4797 0.5278 0.5609 0.6316 0.6947
Full 0.2406 0.5925 0.5925 0.6541 0.6541
PubMed Random 0.3918 0.4420 0.5836 0.6191 0.6191
Ours 0.3974 0.5181 0.7515 0.8052 0.8080
Full 0.3921 0.4484 0.6082 0.6082 0.6670
GAT Cora Random 0.3124 0.3623 0.4603 0.6488 0.7634
Ours 0.4954 0.6007 0.6506 0.7726 0.8299
Full 0.3216 0.3604 0.4713 0.6654 0.7782
Citeseer Random 0.6451 0.6451 0.6451 0.6917 0.6917
Ours 0.5594 0.5504 0.6105 0.7053 0.7504
Full 0.6331 0.6331 0.6617 0.7158 0.7158
PubMed Random 0.4027 0.4603 0.5787 0.5787 0.5787
Ours 0.4174 0.3977 0.4154 0.5580 0.6528
Full 0.3987 0.4415 0.5455 0.5455 0.5455
GraphSAGE Cora Random 0.4824 0.7689 0.7745 0.7893 0.8041
Ours 0.4991 0.7006 0.7486 0.8115 0.8244
Full 0.4880 0.7745 0.7837 0.8059 0.8244
Citeseer Random 0.4797 0.6737 0.6902 0.7038 0.7113
Ours 0.4857 0.6451 0.6692 0.7188 0.7624
Full 0.4812 0.6872 0.6872 0.6992 0.7218
PubMed Random 0.3974 0.6429 0.7406 0.7697 0.7735
Ours 0.3977 0.5945 0.6523 0.7210 0.8078
Full 0.3974 0.6378 0.7484 0.7652 0.7662
4 Improving Hypergraph partitioning
In this section, we propose a robustness-aware hypergraph partitioning framework. We first identify and extract a robust sub-hypergraph consisting of vertices that are empirically less sensitive to perturbations, and perform the partitioning algorithm only on this reduced structure. The resulting partition labels are then propagated to the remaining non-robust vertices by exploiting the underlying graph topology. In this way, our method reduces the effective size of the hypergraph (in terms of the number of vertices) while largely avoiding the degradation in solution quality that typically arises from changes to the hypergraph topology. The main contributions of this paper are summarized as follows:
  • We propose a novel robustness-aware hypergraph partitioning method that leverages adversarial robustness evaluation to distinguish robust nodes from non-robust nodes and treats them differently during the partitioning process.
  • We propose a highly novel label propagation scheme for propagating partition labels within hypergraphs.
  • We conduct extensive experiments on diverse hypergraph instances extracted from realistic VLSI design problems and demonstrate that our approach achieves clearly improved performance over existing hypergraph coarsening methods.
4.1 Method
Our method is built on the key idea of treating nodes that are robust to perturbations separately from those that are vulnerable. For nodes that are less susceptible to perturbations, we directly apply a hypergraph partitioning algorithm to obtain their partitions. In contrast, for nodes that are more easily affected by such perturbations, we do not rely on hypergraph spectral partitioning; instead, we infer their partition membership based on their relationships with the robust nodes and propagate the partitioning results from the robust subset to these vulnerable nodes. Specifically, as shown in Figure 2, our method consists of the following five phases.
Figure 2: Overview of the proposed method.
4.1.1 Phase 1: Hypergraph Spectral Embedding
We first perform a spectral embedding of the hypergraph by mapping it into the feature space spanned by the $K$ eigenvectors of its Laplacian matrix corresponding to the smallest eigenvalue magnitudes.
4.1.2 Phase 2: Manifold Capturing
Data manifolds are commonly modeled using graphs, with k-nearest neighbor (k-NN) graphs being one of the most prevalent choices. In this stage, we construct a k-NN graph to capture the underlying manifold structure of the spectrally embedded hypergraph. A standard k-NN construction procedure has a time complexity of $O(|N|^{2})$, where $|N|$ denotes the number of nodes, which becomes prohibitive for large-scale datasets. To address this issue, we adopt an approximate k-NN graph construction algorithm [25], which extends a probabilistic skip-list–like structure and reduces the complexity to $O(|N|\log|N|)$, making it far more suitable for large-scale data processing.
4.1.3 Phase 3: Adversarial Robustness Evaluation
Let $L_{X}$ and $L_{Y}$ denote the Laplacian matrices associated with the hypergraph $H$ and the $k$-NN graph constructed from the spectrally embedded hypergraph in phase 2, respectively. Following the theoretical framework in [24], we exploit the dominant generalized eigenvalues of $L_{Y}^{+}L_{X}$ and their corresponding eigenvectors to characterize the sensitivity of each sample to noise under a given model. In particular, we compute a weighted eigensubspace matrix $V_{s}\in\mathbb{R}^{|V|\times s}$ to obtain a spectral embedding of the input manifold $G_{X}=(V,E_{X})$, where $|V|$ is the number of nodes. The matrix $V_{s}$ is defined as
$V_{s}=\begin{bmatrix}v_{1}\sqrt{\zeta_{1}},&v_{2}\sqrt{\zeta_{2}},&\ldots,&v_{s}\sqrt{\zeta_{s}}\end{bmatrix}$,
where $\zeta_{1}\geq\zeta_{2}\geq\cdots\geq\zeta_{s}$ are the largest $s$ eigenvalues of $L_{Y}^{+}L_{X}$, and $v_{1},v_{2},\ldots,v_{s}$ are the associated eigenvectors.
Using $V_{s}$, each node $p$ in $G_{X}$ is mapped to an $s$-dimensional representation given by the $p$-th row of $V_{s}$. As described in [24], the stability of an edge $(p,q)\in E_{X}$ can then be measured by the spectral embedding distance between nodes $p$ and $q$:
$\|V_{s}^{\top}e_{p,q}\|_{2}^{2}.$
To quantify the robustness of a node $p$, a score termed the Spade score is introduced as the average embedding distance to its neighbors on the input manifold:
$\text{Spade}(p)=\frac{1}{|N_{X}(p)|}\sum_{q\in N_{X}(p)}\|V_{s}^{\top}e_{p,q}\|_{2}^{2},$
where $N_{X}(p)$ denotes the neighbor set of node $p$ in $G_{X}$. A larger value of $\text{Spade}(p)$ indicates that node $p$ is more vulnerable to adversarial perturbations.
4.2 Phase 4: Sub-Hypergraph Partitioning
For the nodes in the hypergraph $H$ we first sort them in descending order according to their Spade scores and select the bottom portion to form the robust node set. Using these robust nodes together with the hyperedges that connect them in the original hypergraph, we construct a robust sub-hypergraph. We then apply a standard hypergraph partitioning algorithm to this sub-hypergraph to obtain a partition of the robust nodes.
4.3 Label Propagation from Robust to Non-Robust Nodes
The nodes that are not included in the robust node set are referred to as non-robust nodes. Since these nodes are vulnerable to adversarial perturbations and attacks that affect hypergraph spectral partitioning, we do not directly apply the hypergraph spectral partitioning algorithm to determine their partitions. Instead, we infer their partition membership based on their relationships with the robust nodes, effectively propagating the partitioning result from the robust subset to the non-robust nodes. Specifically, for each non-robust node, we identify the robust node with which it co-occurs in hyperedges in $H$ most frequently, and assign the non-robust node to the same partition as that robust node.
The complete algorithm of the proposed method has been shown in Algorithm 1.
Algorithm 1 Robust Hypergraph Partitioning
Input: A hypergraph $H=(V,E)$
Output: Partitioned hypergraph.
1: Extract the adjacency matrix AA from the input hypergraph $H$.
2: Compute the Laplacian matrix $\mathbf{L_{H}}$ for the graph $H$.
3: Compute the eigenvectors $u_{1}$$u_{k}$ that correspond to the bottom k nonzero eigenvalues of AA;
4: Construct $U\in\mathbb{R}^{n\times k}$, with $k$ eigenvectors of AA stored as columns;
5: Construct a $k$-nearest neighbor graph $G=(V,E)$ from $U$.
6: Compute the Laplacian matrix $\mathbf{L_{U}}$ for the graph $U$.
7: Compute the SPADE score of each node.
8: Form a robust node set from nodes with small SPADE scores, while constructing a non-robust node set from nodes with large SPADE scores.
9: Construct a sub-hypergraph $S$ consisting of the robust node set and all hyperedges in $H$ that connect only these robust nodes.
10: For each non-robust node, count the number of hyperedges in $H$ where it appears together with each robust node.
11: Apply standard hypergraph spectral partitioning on $S$.
12: For each non-robust node, identify the robust node with which it co-occurs in hyperedges in $H$ most frequently, and assign the non-robust node to the partition containing that robust node.
13: Return the partitioned hypergraph.
4.4 Experiments
We perform extensive experiments to evaluate the performance of the proposed spectral hypergraph coarsening method on a suite of VLSI-related benchmark datasets, from “ibm01” to “ibm04”Datasets are available at: https://vlsicad.ucsd.edu/UCLAWeb/cheese/ispd98.html.. To ensure a fair comparison, we also apply several existing hypergraph coarsening techniques to the same set of benchmarks and compare their results against ours. For brevity, we refer to the star and clique models as S and C, respectively. The experimental results are shown in Table 3. It can be seen that our method consistently outperforms all competing methods across all four datasets.
Table 3: The average local conductance ($\Phi$) comparison between our methods and other methods
Hypergraph V E Spectral S Spectral C Metis S Metis C hMETIS HyperSF Ours
ibm01 12752 14111 0.94 0.81 0.84 0.78 0.65 0.44 0.28
ibm02 19601 19584 0.94 0.84 0.83 0.79 0.69 0.52 0.43
ibm03 23136 27401 0.95 0.86 0.85 0.79 0.67 0.48 0.41
ibm04 27507 31970 0.95 0.84 0.84 0.78 0.68 0.47 0.40
5 Strengthen GNN Defenses via Adversarial Robustness Evaluation
The proposed method for improving the robustness of GNNs, which consists of five key phases. Phase (1) learns a latent node representation: we train a GNN and use the activations of its hidden layer as embeddings for the original feature matrix. In Phase (2), we construct $k$-NN graphs to capture the underlying manifold structure of the embedded feature space. In Phase (3), we quantify the robustness of each node by jointly leveraging information from the input graph and the learned output manifold. In Phase (4), we select a subset of robust nodes, extract the induced subgraph consisting only of them from the original graph, and apply a GNN to classify these robust nodes. In Phase (5), we perform label propagation based on the connectivity patterns between robust and non-robust nodes. In the rest of this section, we provide a detailed description of these five phases.
5.1 Method
5.1.1 Phase 1: Learning GNN Latent Representations
[24] introduced a framework that analyzes the adversarial robustness of individual edges by constructing a bijective distance mapping between the manifolds of a model’s input space and output space.
In our setting, we first train a GNN on the original feature matrix together with the given input graph to obtain a well-fitted model. After training, this GNN is used to map the input features into a latent representation space. Since the last layer of a GNN is typically tailored to the downstream task (for example, a softmax layer in the case of multi-class node classification), we use the activations of the penultimate hidden layer as the embedding of each node, rather than the final task-specific outputs.
5.1.2 Phase 2: Capturing the Data Manifold
A common way to approximate the underlying data manifold is to represent it with a graph, where nodes correspond to data points and edges reflect local neighborhood relationships. Among such constructions, $k$-nearest neighbor ($k$-NN) graphs are particularly popular. In this phase, we construct a $k$-NN graph on top of the embedded feature matrix obtained from the GNN, so as to model the manifold structure in the latent space.
The standard $k$-NN graph construction procedure has a time complexity of $O(|N|^{2})$, where $|N|$ denotes the number of nodes, which becomes prohibitive for large-scale datasets. To address this scalability issue, we adopt an approximate $k$-NN graph construction method [25]. This approach builds on an extension of probabilistic skip-list–like structures to significantly reduce the computational cost, achieving a time complexity of $O(|N|\log|N|)$ and thus making it suitable for large data regimes.
5.1.3 Phase 3: Spectral Adversarial Robustness Evaluation
Let $L_{X}$ and $L_{Y}$ denote the Laplacian matrices of the original input graph and the $k$-NN graph constructed in the embedded feature space, respectively. Following the theoretical framework in [24], we leverage the leading generalized eigenvalues of $L_{Y}^{+}L_{X}$ together with their associated eigenvectors to estimate how sensitive each sample is to noise under a specified model. Specifically, we form a weighted eigensubspace matrix $V_{s}\in\mathbb{R}^{|V|\times s}$, which serves as a spectral embedding of the input manifold $G_{X}=(V,E_{X})$, where $|V|$ denotes the number of nodes. The matrix $V_{s}$ is defined as
$V_{s}=\begin{bmatrix}v_{1}\sqrt{\zeta_{1}},&v_{2}\sqrt{\zeta_{2}},&\ldots,&v_{s}\sqrt{\zeta_{s}}\end{bmatrix}$,
with $\zeta_{1}\geq\zeta_{2}\geq\cdots\geq\zeta_{s}$ being the top $s$ eigenvalues of $L_{Y}^{+}L_{X}$, and $v_{1},v_{2},\ldots,v_{s}$ the corresponding eigenvectors.
Under this construction, each node $p$ in $G_{X}$ is represented in $\mathbb{R}^{s}$ by the $p$-th row of $V_{s}$. As stated in [24], the stability of an edge $(p,q)\in E_{X}$ can be assessed via the distance between nodes $p$ and $q$ in the spectral embedding:
$\|V_{s}^{\top}e_{p,q}\|_{2}^{2}.$
To summarize node-level robustness, we adopt the Spade score, defined as the mean embedding distance from node $p$ to its neighbors on the input manifold:
$\text{Spade}(p)=\frac{1}{|N_{X}(p)|}\sum_{q\in N_{X}(p)}\|V_{s}^{\top}e_{p,q}\|_{2}^{2},$
where $N_{X}(p)$ is the neighbor set of node $p$ in $G_{X}$. A higher $\text{Spade}(p)$ suggests that node $p$ is more susceptible to adversarial perturbations.
5.1.4 Phase 4: Sub-graph Classification
For the nodes in the original graph, we first rank them in descending order by their Spade scores and then take the lowest-scoring fraction to define a robust node set. Using these robust nodes and the edges among them inherited from the original graph, we build a robust subgraph. Subsequently, we train a GNN on this robust subgraph to perform node classification for the robust nodes.
5.1.5 Phase 5: Label Propagation from Robust to Non-Robust Nodes
The nodes that are not included in the robust node set are referred to as non-robust nodes. Because these nodes are more susceptible to adversarial perturbations and attacks that can compromise GNN predictions, we do not apply the GNN to classify them directly. Instead, we infer their class labels from their associations with the robust nodes, i.e., we propagate the predictions obtained on the robust subset to the non-robust nodes. Concretely, for each non-robust node, we find the robust node with which it shares the largest number of connections, and assign the non-robust node to the same class as that robust node.
5.2 Experiments
We evaluate robustness using an independent (non-cumulative) white-box greedy edge-deletion attack. For each target test node $v$, we start from a fresh copy of the clean graph $G$ and perform at most $T$ deletion steps. At step $t$, we enumerate all incident edges $(v,u)$, temporarily remove each candidate edge, and run a forward pass to measure the logit of the true class $y_{v}$ at node $v$. We then permanently delete the edge that yields the largest decrease in this true-class logit. The attack stops early if $v$ becomes misclassified, if no incident edges remain, or if removing any remaining edge no longer decreases the true-class logit. This procedure is applied independently to each target node (i.e., perturbations are not shared across targets). We select as effective targets those test nodes that are correctly classified on the clean graph by both the baseline and our method, and report the post-attack accuracy on these targets (equivalently, $1-\mathrm{ASR}$). The experimental results are shown in Table 4. It can be seen that our method is very effective.
Table 4: Post-attack accuracy on effective target nodes under the independent white-box greedy edge-deletion attack (Cora).
Model Effective Targets Baseline Acc. Ours Acc.
GAT 200 0.7450 0.7900
GCN 200 0.7300 0.7650
6 Conclusion
This paper adopts spectral adversarial robustness evaluation to identify and select the most robust nodes from a given graph. Building on this selection, we propose three algorithms that respectively improve the training quality of GNNs, enhance the performance of hypergraph partitioning, and strengthen the adversarial robustness of GNNs against attacks. Extensive experiments demonstrate the effectiveness of all three algorithms.